OpenStudy (anonymous):

Consider a Hydrogen atom with the electron in the n = 7 shell. What is the energy of this system? (The magnitude of the ground state energy of the Hydrogen atom is 13.6 eV.)

OpenStudy (amriju):

see rydberg eqn

OpenStudy (amriju):

(1/n^2-1/n'^2).*Rm=1/l here n=1 n'=7 Rm is the rydberg constant.. l= wavelength once l is known..find f( frequency) [use f=c/l] hf , h= plancks const.. will give u energy.. I am looking for the value of Rydbergs constant in the net..i dont remember it

OpenStudy (amriju):

ok i saw it...Rm=Rinf(1-m/M), m is the mass of electron, M is the mass of atom hydrogen. now R inf is another constant...how do we find it...? look at my previous eqn...put n'=infinity so than 1/n'=0. and put n=1and replace Rm by Rinf so Rinf=1/l. Now this value of l gives u the wavelength emitted when the electrons jumps from state 1 to infinity..in other words E=h*c/l in this case will give u the energy at the ground state. E has been provided to u..find l and then Rinf..subsequently Rm..

OpenStudy (amriju):

then use the eqn to find energy

OpenStudy (anonymous):

okay i am a little confused on how to find Rinf

OpenStudy (amriju):

when i explain i use terms like baby n all...please dont mind...

OpenStudy (amriju):

look dear...wthe rydbergs eqn u see calculates the energy released or gained when an electron of a 1 electron systems jumps from one state to another.. can u understand this much?

OpenStudy (amriju):

@witthan1

OpenStudy (amriju):

can u understand that much...can i proceed?

OpenStudy (anonymous):

sure

OpenStudy (amriju):

now realise one thing...electrons have negative energy..which is why they are bound..if they had a positive finite energy..they would move away is it not? So, understand that at distance infinity the energy of an electron due to the atom is 0. Can I say now that if I take an electron with energy -E at the ground state to infinity, it will absorb E amount of Energy, so that its energy becomes 0 at infinity? So can I say that if I use Rydbergs eqn to find energy( it gives u wavelength U can calculate energy from there) absorbed by an electron in moving from ground state to infinity u would have the magnitude of the Energy of the elctron at ground state? Do u understand sweetheart?

OpenStudy (amriju):

hey..

OpenStudy (anonymous):

gotcha thanks

OpenStudy (amriju):

:) wc